OpenStudy (anonymous):

The function y=-0.003x^2+1.60x models the path of a kicked soccer ball. The height is y, the distance is X, and the units are meters. How far does the soccer ball travel? How high does the soccer ball go? describe a reasonable domain and range for the function.

OpenStudy (anonymous):

How much will the height (y) be when the ball falls down?

OpenStudy (anonymous):

i dont know..:/ if its down, then 0?

OpenStudy (anonymous):

Exactly! So for what distance x do you have that -0.003x^2+1.60x = 0?

OpenStudy (anonymous):

not sure..:/ what would it be?

OpenStudy (anonymous):

can you please help me? what do i do first?

OpenStudy (anonymous):

\[ -0.003x^2+1.60x = 0\]\[x( -0.003x+1.60) = 0\]\[x=0\] or \[-0.003x+1.60 = 0\] Can you solve \[-0.003x+1.60 = 0\]?

OpenStudy (anonymous):

how do i solve it? what do i need to do?

OpenStudy (anonymous):

Well, you can isolate the x, so you get \[-0.003x=-1.60 \rightarrow x = 1.60/0.003 = 533.333\]

OpenStudy (anonymous):

To find how high the ball goes you have to find the vertex of the parabola, do you know how to do that?

OpenStudy (anonymous):

no